OpenStudy (anonymous):

72 is what percent of 480? Write and solve an equation.

OpenStudy (shamil98):

\[\huge \frac{ part }{ whole } = \frac{ percent }{ 100 }\]

OpenStudy (anonymous):

480 / 72 = 100 / x x = 7200 / 480

OpenStudy (shamil98):

\[\huge \frac{ 72 }{ 480 } = \frac{ percent }{ 100 }\] solve for the percent. , cross multiply.

OpenStudy (anonymous):

do a problem you know first 72 is what percent of 480? Write and solve an equation. 72 100 you know it is 72 right :) so you know you got that by 72/100 then shift decimal 2 places to the right 72/480 then shift

OpenStudy (perl):